Tags:
create new tag
view all tags

20120521 Shift

Shift Plan

  • Swap to LW optics.
  • OTR polariser and quick quad scan - keep quick - will do proper measurement later
  • Background tuning - to ~ 6000 counts at 10dB
  • LW spatial alignment.
  • Timing.
  • Collisions.

Scans

  • Initial Vertical
  • Initial Phase
  • Initial Horizontal

  • 2000 pulse background log - laser off
  • 2000 pulse log - on peak of collisions
  • Laser energy correlation
  • Charge correlation

  • Dispersion measurement / correction - vertical before & after
  • Coupling correction scan - 5 - 7 vertical scans - set to minimum and redo vertical
  • Long range horizontal scan
  • Telescope scan - repeated vertical - set to minimum - redo vertical

  • QM14FF quad scan
  • Switch to OTR to measure quad scan at exact same positions!

From this point on - if time allows:

  • Vertical slices at different horizontal locations - around 7 - 9
  • LW scan of beam, then after carbon wire at IP, the LW on peak and wire scan repeat

Shift

17:00

Shift starts. Set file from previous shift set12may21_1703.dat

Access to reset energy meter.

Beam on 17:20.

Previous beam tuning shift using set12may18_1200.dat which are from our previous shift, so beam tuning shift running our optics anyway. Load set 12may18_1733.dat which was the setting file just after the earthquake last shift. This was at the end of a quad scan so QM14FF is -94.260 - reset to -88.260 which was the middle of the quad scan before.

Beam tuning for background - 4500 counts with 10dB attenuation

17:45

Beam off, access to check cables - no communication with chamber computer.

18:15

  • Beam back on. Oil leak near extraction kicker - took 0.5hrs to fix.

19:35

  • OTR Polariser scan. Minimum rotation angle 4 degrees for analysis. Minimum polariser angle 12.9 degrees.
  • OTR minimum size from polariser scan 1.36 +- 0.02

E-beam Alignment

During OTR analysis, perform LW alignment scans.

  • Laser referenced earlier to 9445 Manipulator Vertical
  • Chamber vertical - set to 820 - only 10 microns different from last time
  • Chamber horizontal - set vetical to 550 - horizontal minimum at 500

  • Set file 12may21_1951.dat after tuning for low background (around 5k at 10dB)

Horizontal Alignment Vertical Alignment
20120521_1934_lws_xy_errbar_plot.png 20120521_1919_lws_xy_errbar_plot.png

Timing

  • Retired to the laser hut to do timing
  • Much trouble with phase wrapping around 0V but optimum being around 0.5V
  • Tried various lengths of cable addition to TD2 for regen but not much difference
  • Regen performance dropping as laser shifting with time
  • Eventually after retuning laser, found peak at 7.0V. Phase scan quite noisy though

Collisions

Found collisions at 850um chaver, but amplitude very low - 1000 on top of 5000. Phase was correct as was horizontal and vertical alignment. Background has slightly increased but still around 6000 so not too bad. Laser energy low 12mJ vs 16mJ last week (measured post ip with attenuator). Using attenuation factor, this corresponds to roughly 70mJ now vs 100mJ previously. Note the beam is slightly clipped on the energy meter and the mirrors are lossy. However, the energy meter is highly linear with the laser pulse energy.

Vertical Scan RMS Plot
20120521_2059_lws_xy_errbar_plot.png 20120521_2059_lws_rms.png

Timing & Collisions

  • Collisions were very low amplitude - retuned regen. Then had to redo timing.
  • Much trouble with timing

Added small amount of cable to clock line for TD2 for regen timing. However, whilst this shifted the phase shifter voltage that the overlap was found, the overlap was still at a wrapping point where it was unstable. Tried several different cables and moved the overlap voltage from 0V, to 3V, to 5V, 7V or thereabouts. In each case still wrapping. Alex had the idea of adding a little bit of cable to the Qswitch input trigger in the top patch panel. This worked quite well. Stable overlap.

20120521_2245_lws_xy_fit_plot.png

Retiming Phase Scan

Initial Vertical Corresponding Nonlinear Vertical Scan
20120521_2252_lws_xy_errbar_plot1.png 20120521_2313_lws_xy_errbar_plot.png

Changed attenuation to 6dB (from 10dB) to take advantage of the dynamic range of the ADC.

Correlation Logs

  • 1000 pulse log on peak 20120521_2318_log.dat
  • 1000 pulse log with no laser and displaced 20120521_2334_log.dat

00:10

  • On peak signal very jumpy - timing unstable.
  • Added further bit of cable to QSwitch - much better. Phase Scan to find collisions. Set to 6.6V

  • lwLog> 20120522_0016_log npulse=1000.0 - On collisions
  • lwScan> 20120522_0030_lws axis=chaver, npulse=5.0, start=742.302428635, stop=942.302428635, nstep=31.0 - Telescope at 11
  • lwScan> 20120522_0036_lws axis=chaver, npulse=5.0, start=750.0, stop=950.0, nstep=31.0 - Telescope 10
  • lwScan> 20120522_0039_lws axis=chaver, npulse=5.0, start=745.0, stop=945.0, nstep=101.0 - Telescope 9
  • lwScan> 20120522_0051_lws telescope 8 rms fit 16.4 um
  • lwScan> 20120522_0058_lws telescope 7 rms fit 17.2 um
  • lwScan> 20120522_0112_lws telescope 6 rms fit 17.5 um

Timing unstable again. Phase scan, new peak at 7.6V.

  • lwScan> 20120522_0129_lws telescope 6 rms fit 26.3 um
  • lwScan> 20120522_0136_lws telescope 11 rms fir 18.0 um
  • lwScan> 20120522_0141_lws telescope 12 rms fit 15.3 um

Timing unstable. Script can't fit phase scan (wScan> 20120522_0147_lws).

02:00

Chamber not centred horizontally. Moved to increase lw signal, attenuation increased to 8dB.

  • lwScan> 20120522_0226_lws CV: 850, CH: -400 - 800, 120 steps, n = 3, telescope 12. Scan didn't complete.

Change to DAQ timing - lwScan> 20120522_0251_log shows good correlation between Cherenkov signal and charge.

Synchronicity... Again

So linac was playing up and there were major jumps in charge as Alex tried to fix this. This showed that in the EDM that the REFC1:amp and our cherenkov signal were exactly 1 pulse out with the laser-wire pvs leading by one pulse. LW dB is much simpler than CBPM so publishes quicker. I tried triggering from REFC1:amp, but this didn't resolve the problem. I wrote a piece of python to delay our software trigger by 100ms and this improved the correlations significantly. In the end Stewart informed me that the REFC1:amp is one of the first to be published as it's hardware based where as the positions etc are published up to 200ms later due to processing. Therefore, the LW DAQ now triggers from cbpm:xpos which is the last to be published of the cbpm pvs.

Charge correlations are significantly better!

20120522_0251_log_correlation_plot.png

Background Correlation - No Laser and displaced from peak

03:05

Horizontal scan again.

20120522_0320_lws_xy_errbar_plot.png

Vertical Scan After Horizontal Centering

vertical scan 20120522_0320_lws rms fit 9um centre 847.

vertical scan 20120522_0332_lws rms fit 9.3um.

Timing scan 20120522_0420_lws, centre at phase 7.6V.

Telescope Scan

Telescope Position Filename RMS Integral Sigma
6mm 20120522_0425_lws 10.7 2.87 +/- 0.19
7mm 20120522_0432_lws 10.1 3.27 +/- 0.16
8mm 20120522_0439_lws 8.5 2.05 +/- 0.15
9mm 20120522_0446_lws 9.1 2.29 +/- 0.18
11mm 20120522_0455_lws 8.3 1.91 +/- 0.20
13mm 20120522_0506_lws 8.5 1.44 +/- 0.23
15mm 20120522_0512_lws 9.3 0.82 +/- 0.29
17mm 20120522_0519_lws 9.4 2.03 +/- 0.22
19mm 20120522_0526_lws 9.7 1.08 +/- 0.26

Minimum of a parabolic fit to the rms data is 12.7mm. Set telescope to 12.7mm and scan phase: 20120522_0535_lws. Max. at 7.58V.

05:45 Dispersion Measurement & Correction

Set file before dispersion correction: 12may22_0546.dat

Perform dispersion measurement: program suggests changing QF1X, QF6X. Set correction as suggested, new set file: 12may22_0552.

Before After
εx = 5.788 +/- 0.60 εx = 5.630 +/- 0.44
εx' = 5.464 +/- 0.33 εx' = 2.973 +/- 0.27
εy = 3.200 +/- 0.34 εy = 3.170 +/- 0.30
εy' = -2.435 +/- 0.25 εy' = -2.454 +/- 0.22

New vertical scan after dispersion: 20120522_0548_lws rms fit 11.1um.

06:05 Coupling Scan

Use QK4X to scan coupling and perform vertical scans at each setting. Original setting 3.090A.

QK4X Current (A) Filename RMS Integral Sigma
3.090 20120522_0548_lws 11.1 0.87 +/ 0.27
6.090 20120522_0557_lws 9.2 0.91 +/ 0.28
9.090 20120522_0604_lws 11.0 0.75 +/ 0.29
12.090 20120522_0611_lws 10.1 0.91 +/ 0.18
20.000 20120522_0620_lws 12.6 0.68 +/ 0.29

Not much difference, set back to 3.090A. Try different settings of QK3X and look at signal peak, not much difference. QK2X seems to make more difference so scan that magnet. Original setting -0.645A.

QK2X Current (A) Filename RMS Integral Sigma
2.000 20120522_0641_lws 12.9 1.52 +/ 0.32
1.500 20120522_0649_lws 13.4 0.61 +/ 0.34
0.800 20120522_0655_lws 11.4 1.63 +/ 0.24
0.000 20120522_0702_lws 12.9 1.89 +/ 0.20
-0.645 20120522_0548_lws 11.1 0.87 +/ 0.27
-1.400 20120522_0709_lws 10.5 1.57 +/ 0.29

QK2X set back to -0.645A.

07:15 Quad scan

QM14FF original setting -88.260A (minimum of OTR scan at beginning of shift).

QM14FF Current (A) Filename RMS Integral Sigma
-88.26 20120522_0548_lws 11.1 0.87 +/ 0.27
-86.26 20120522_0716_lws 9.9 2.44 +/ 0.20
-84.26 20120522_0728_lws 11.5 3.35 +/ 0.36
-82.26 20120522_0736_lws 11.8 3.03 +/ 0.18
-80.26 20120522_0743_lws 11.4 5.15 +/ 0.30
-78.26 20120522_0751_lws 12.3 4.83 +/ 0.35
-90.26 20120522_0802_lws 12.6 2.66 +/ 0.30
-92.26 20120522_0809_lws 13.2 3.48 +/ 0.35
-94.26 20120522_0817_lws 11.9 3.95 +/ 0.38
-96.26 20120522_0823_lws 13.5 3.92 +/ 0.27
-98.26 20120522_0830_lws 14.7 4.57 +/ 0.40

QM14FF back to -88.26A.

Detailed vertical scan 20120522_0840_lws rms fit 20.2, not well centred. Vertical scan 20120522_0849_lws rms fit 14.7.

Set file 12may22_0904.dat

09:00 OTR quad scan

09:30 end of shift

NB OTR quad scan data

Analysed 24/5/2012 LC - rotation angle 3.69 degrees, minimum of scan at -87.718A, sy = 1.748 +/- 0.023um.

  • Smallest fit from Telescope Scan:
    20120522_0512_lws_xy_fit_plot.png
Topic attachments
I Attachment History Action Size Date Who Comment
PNGpng 20120521_1919_lws_xy_errbar_plot.png r1 manage 22.0 K 21 May 2012 - 11:54 LaurieNevay Chamber Vertical Scan
PNGpng 20120521_1934_lws_xy_errbar_plot.png r1 manage 24.2 K 21 May 2012 - 11:53 LaurieNevay Horizontal Chamber Scan
PNGpng 20120521_2059_lws_rms.png r1 manage 39.8 K 21 May 2012 - 13:28 LaurieNevay First Vertical RMS
PNGpng 20120521_2059_lws_xy_errbar_plot.png r1 manage 28.6 K 21 May 2012 - 13:29 LaurieNevay First Vertical Scan
PNGpng 20120521_2245_lws_xy_fit_plot.png r1 manage 33.8 K 21 May 2012 - 15:26 LaurieNevay Retiming Phase Scan
PNGpng 20120521_2252_lws_xy_errbar_plot1.png r1 manage 28.2 K 21 May 2012 - 15:27 LaurieNevay Initial Vertical
PNGpng 20120521_2313_lws_xy_errbar_plot.png r1 manage 27.3 K 21 May 2012 - 15:28 LaurieNevay Nonlinear Vertical Scan
PNGpng 20120522_0251_log_correlation_plot.png r1 manage 58.4 K 21 May 2012 - 19:15 LaurieNevay Background Correlation
PNGpng 20120522_0320_lws_xy_errbar_plot.png r1 manage 32.8 K 21 May 2012 - 21:18 LaurieNevay Vertical Scan After Horizontal Centering
PNGpng 20120522_0512_lws_xy_fit_plot.png r1 manage 46.9 K 21 May 2012 - 22:01 LaurieNevay Smallest fit from Telescope Scan
Edit | Attach | Watch | Print version | History: r15 < r14 < r13 < r12 < r11 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r15 - 23 May 2012 - LauraCorner

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback